when rolling a pair of dice, what is the probability of rolling exactly 3 sums of 6 in exactly 8 rolls?
Answer by oscargut(2103)

P(sum=6) = 5/36

Let S =the number of sums of 6 in 8 rolls
S has a binomial distribution with n= 8 and p = 5/36
P(S =3) = C(8,3)(5/36)*3(31/36)*5 = 0.071
Answer: 0.071
You can ask me more at: mthman@gmail.com
Thanks